eviant:

PLS HELP!!! If f(x) = 5x – 6, what is the equation for f–1(x)?

eviant:

@Vocaloid

dude:

Uh basically find the inverse of the equation, do you know how to do that?

eviant:

no

dude:

\(f(x) = 5x – 6\) Change f(x) with y to make it easier to look at \(y=5x-6\) Now flip x and y and solve for y \(x=5y-6\)

eviant:

I think you add 6 or 5 next to both sides?

dude:

Add 6, the 5 is being multiplied by y

eviant:

x=11

eviant:

x=55

dude:

Wait what, you're not supposed to get an answer like that Add 6 on both sides \(x+6=5y\) Then divide by 5 on both sides \(\large \frac{x+6}{5}=y\) That is your equation

eviant:

oh

eviant:

thanks